OpenStudy (anonymous):

solve -1/3x + y = -1 y=4 +1/3x

OpenStudy (anonymous):

substitute y from the second equation in the first

ganeshie8 (ganeshie8):

\(\huge -\frac{1}{3}x + y = -1 \\\huge y = 4 +\frac{1}{3}x\)

ganeshie8 (ganeshie8):

are these ur equations ?

OpenStudy (anonymous):

add the two of your equations

OpenStudy (anonymous):

and then substitute

ganeshie8 (ganeshie8):

@kmstevens10 hello.. ?

OpenStudy (anonymous):

yes